OpenStudy (anonymous):

Polynomials and Polynomial Functions Unit Test lesson 11 unit 5 HELP PLEASE!

OpenStudy (anonymous):

1. Which equation is most likely represented by the graph above? (1 point)y = (2x – 1)(2x – 3)(2x + 5) y = (2x + 1)(2x + 3)(2x – 5) y = –(2x + 1)(2x + 3)(2x – 5) y = (2x + 1)(2x + 3)(2x + 5) 2. The graph of is shown below. What are the apparent zeros of the function graphed above? (1 point){–1, 2.5} {–17, 5} {–4, 0, 2} {–2, 0, 4} 3. What is a polynomial function in standard form with zeros 1, 2, –2, and –3? (1 point)x4 + 2x3 + 7x2 – 8x + 12 x4 + 2x3 – 7x2 – 8x + 12 x4 + 2x3 – 7x2 + 8x + 12 x4 + 2x3 + 7x2 + 8x + 12 Multiple Choice 4. What are the zeros of the function? What are their multiplicities? f(x) = 5x3 – 5x2 – 30x (1 point)The numbers 3, –2, and 0 are zeros of multiplicity 1. The numbers 3, –2, and 0 are zeros of multiplicity 2. The numbers –3, 2, and 0 are zeros of multiplicity 1. The numbers –3, 2, and 0 are zeros of multiplicity 2. 5. What are the real or imaginary solutions of the polynomial equation? x3 – 8 =0 (1 point)1 + isquare root 3, and 1 – isquare root 3 2, –1 + isquare root 3, and –1 – isquare root 3 2, 1 + 2isquare root 3, and 1 – 2i square root 3 2, 2 +2isquare root 3, and 2 – 2isquare root 3 6. Which is equivalent to the following expression? (3m2 + 2mn – n2) + (m2 + 4mn – n2) (1 point)4m2 + 6mn – 2n2 4m4 – 6mn – 2n4 3m2 – 6mn + 2n2 4m2 + 6mn + 2n2 7. Which is a factored form of 27x3 + 64? (1 point)(3x + 4)(9x2 – 24x + 16) (3x + 4)(9x2 + 24x + 16) (3x + 4)(9x2 – 12x + 16) (3x + 4)(9x2 + 12x + 16) Multiple Choice 8. Divide –2x3 – 4x2 + 3x + 2 by x – 3. (1 point)–2x2 – 10x – 27, R –79 –2x2 – 10x –27 –2x2 – 10x – 27, R 83 –2x2 – 10x + 27, R 83 9. Find the roots of the polynomial equation. x3 – x2 + x + 39 (1 point)1 + 3i, 1 – 3i, –3 1 + 3i, 1 – 3i, 3 –2 + 3i, –2 – 3i, –3 2 + 3i, 2 – 3i, –3 10. Which correctly describes the roots of the following cubic equation? x3 + 6x2 + 11x + 6 (1 point)One real root, two complex roots Two real roots and one complex root Three real roots, two of which are equal in value Three real roots, each with a different value 11. What is the second term of (p + q)5? (1 point)25p4 5p4q 25p4q 125p4q 12. One zero of f(x) = x3 + 2x2 – 5x – 6 is 2. What are other zeros of the function? (1 point)1 and 3 –1 and –3 2 and –3 2 and 3 13. Which of the following quartic functions has x = –2 and x = –3 as its only two real zeros? (1 point)x4 – 5x3 + 7x2 – 5x + 6 x4 + 5x3 + 7x2 + 5x + 6 x4 + 5x3 + 7x2 – 5x – 6 x4 – 5x3 + 7x2 + 5x – 6 14. Work Pad Note: Remember to show all of the steps that you use to solve the problem. You can use the comments field to explain your work. Your teacher will review each step of your responses to questions 13–19 to ensure you receive proper credit for your answers. The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 times the height, what should each dimension be? Hint: Let x represent one of the dimensions, and then define the other dimensions in terms of x. (2 points) 15. What are the real or imaginary solutions of the polynomial equation? x4 – 41x2 = –400 (2 points) 16. Use synthetic division to find P(–3) for P(x) = x4 – 2x3 – 4x + 4. (2 points) 17. Find all the zeroes of the equation. x4 – 6x2 – 7x – 6 = 0 (2 points) 18. Use Pascal’s triangle to expand the binomial. (d – 5y)6 (2 points) 19. What is the equation of y = x3 with the given transformations? vertical compression by a factor of 1 over 7, horizontal shift 8 units to the left, reflection across the x-axis (2 points) 20. What are all the real zeroes of y = (x –12)3 – 7? (2 points)

OpenStudy (anonymous):

don't worry about work pad

OpenStudy (anonymous):

Hey I sent you a message. Let me know if you get it!

OpenStudy (anonymous):

did you get the answers?

OpenStudy (anonymous):

He/ She gave me all of the wrong answers. I wouldn't recommend taking them.

OpenStudy (anonymous):

With all due respect, I am not the type of person to sabotage someones test. It might be possible that your test was in a different order

OpenStudy (anonymous):

Hey i have the same one can i have the answers

OpenStudy (anonymous):

:D p.s. this is a potato

OpenStudy (hackergod):

me too!

OpenStudy (anonymous):

Same here! XD

OpenStudy (janewelcomes):

I fourth that!

OpenStudy (anonymous):

Ahaha XD check your email @JaneWelcomes

OpenStudy (anonymous):

1. Which equation is most likely represented by the graph above? (1 point) Correct answer (1 pt) y = (2x – 1)(2x – 3)(2x + 5) Unselected answer (0 pts) y = (2x + 1)(2x + 3)(2x – 5) Unselected answer (0 pts) y = –(2x + 1)(2x + 3)(2x – 5) Unselected answer (0 pts) y = (2x + 1)(2x + 3)(2x + 5) 1 /1 point 2. Shown below is the graph of . What are the apparent zeros of the function graphed above? (1 point) Unselected answer (0 pts) {–0.7, –2.7} Unselected answer (0 pts) {–10.2, 10.2} Unselected answer (1 pt) {–2, 1, 4} Incorrect Answer (0 pts) {–4, –1, 2} 0 /1 point 3. What is a polynomial function in standard form with zeros 1, 2, –2, and –3? (1 point) Unselected answer (0 pts) x4 + 2x3 + 7x2 – 8x + 12 Correct answer (1 pt) x4 + 2x3 – 7x2 – 8x + 12 Unselected answer (0 pts) x4 + 2x3 – 7x2 + 8x + 12 Unselected answer (0 pts) x4 + 2x3 + 7x2 + 8x + 12 1 /1 point Multiple Choice 4. What are the zeroes of the function? What are their multiplicities? f (x) = 4x3 – 20x2 + 24x (1 point) Incorrect Answer (0 pts) The numbers –3, –2, and 0 are zeroes of multiplicity 1. Unselected answer (0 pts) The numbers 3, 2, and 0 are zeroes of multiplicity 2. Unselected answer (0 pts) The numbers –3, –2, and 0 are zeroes of multiplicity 2. Unselected answer (1 pt) The numbers 3, 2, and 0 are zeroes of multiplicity 1. 0 /1 point 5. What are the real or imaginary solutions of the polynomial equation?x3 – 64 = 0 (1 point) Unselected answer (0 pts) 4, –1 + , –1 + Incorrect Answer (0 pts) 4, 1 + , 1 + Unselected answer (1 pt) 4, –2 + 2i3, –2 – 2i3 Unselected answer (0 pts) 4, –2 + , 2 + 0 /1 point 6. Which is equivalent to the following expression? (1 point) Incorrect Answer (0 pts) 3p2 + 8pq – 3q2 Unselected answer (0 pts) 4p2 + 8pq + q2 Unselected answer (0 pts) 4p4 + 8pq – 3q4 Unselected answer (1 pt) 4p2 + 8pq – 3q2 0 /1 point 7. Which is a factored form of 8x3 + 27? (1 point) Unselected answer (0 pts) (2x + 3)(2x + 3)(2x + 3) Unselected answer (1 pt) (2x + 3)(4x2 – 6x + 9) Incorrect Answer (0 pts) (2x – 3)(4x2 + 6x + 9) Unselected answer (0 pts) (2x – 9)(4x2 + 18x +81) 0 /1 point Multiple Choice 8. Divide –3x3 – 4x2 + 4x +3 by x – 2. (1 point) Incorrect Answer (0 pts) –3x2 + 2x + 24 Unselected answer (1 pt) –3x2 – 10x – 16, R – 29 Unselected answer (0 pts) –3x2 – 10x – 16 Unselected answer (0 pts) –3x2 + 2x + 24, R 35 0 /1 point 9. Find the roots of the polynomial equation. x3 – 4x2 + x + 26 (1 point) Unselected answer (1 pt) 3 + 2i, 3 – 2i, –2 Unselected answer (0 pts) 3 + 2i, 3 – 2i, 2 Unselected answer (0 pts) –3 + 2i, –3 – 2i, 2 Incorrect Answer (0 pts) –3 + 2i, –3 – 2i, –2 0 /1 point 10. Which correctly describes the roots of the following cubic equation? x3 – 5x2 + 3x + 9 = 0 (1 point) Unselected answer (0 pts) Three real roots, each with a different value Unselected answer (0 pts) One real root and two complex roots Unselected answer (1 pt) Three real roots, two of which are equal in value Incorrect Answer (0 pts) Two real roots and one complex root 0 /1 point 11. What is the second term of (s + v)5? (1 point) Unselected answer (0 pts) 25s4 Incorrect Answer (0 pts) 25s4v Unselected answer (1 pt) 5s4v Unselected answer (0 pts) 125s4v 0 /1 point 12. One zero of f(x) = x3 – 2x2 – 5x + 6 is 1. What are other zeros of the function? (1 point) Unselected answer (1 pt) 3 and –2 Incorrect Answer (0 pts) –3 and 2 Unselected answer (0 pts) 1 and 2 Unselected answer (0 pts) 1 and –3 0 /1 point 13. Which of the following quartic functions has x = –2 and x = –3 as its only two real zeros? (1 point) Unselected answer (0 pts) x4 – 5x3 + 7x2 – 5x + 6 Correct answer (1 pt) x4 + 5x3 + 7x2 + 5x + 6 Unselected answer (0 pts) x4 + 5x3 + 7x2 – 5x – 6 Unselected answer (0 pts) x4 – 5x3 + 7x2 + 5x – 6

OpenStudy (anonymous):

I could kiss you!

OpenStudy (anonymous):

hahaha :)

OpenStudy (snowcrystal):

you have totally saved my neck right here :) thank you!!!!!